# ConTextTab: A Semantics-Aware Tabular In-Context Learner

### v2

@@ Abstract
Tabular in-context learning (ICL) has recently achieved state-of-the-art (SOTA) performance on several tabular prediction tasks. Previously restricted to classification problems on small tables, recent advances such as TabPFN [17] and TabICL [28] have extended its use to larger datasets. Although current table-native ICL architectures are architecturally efficient and well-adapted to tabular data structures, their exclusive training on synthetic data limits their ability to fully leverage the rich semantics and world knowledge contained in real-world tabular data. At the other end of the spectrum, tabular ICL models based on pretrained large language models such as TabuLa-8B [11] integrate deep semantic understanding and world knowledge but are only able to make use of a small amount of context due to inherent architectural limitations. With the aim to combine the best of both these worlds, we introduce ConTextTab, integrating semantic understanding and alignment into a table-native ICL framework. By employing specialized embeddings for different data modalities and by training on large-scale real-world tabular data, our model is competitive with SOTA across a broad set of benchmarks while setting a new standard on the semantically rich CARTE benchmark. Code and checkpoints are available at: https://github.com/SAP-samples/contexttab.
@@ 1 Introduction
Tables with information spread across rows and columns remain a predominant data format in many real-world applications [3], making their understanding through machine learning algorithms critical. Despite the great success of deep learning approaches in natural language processing and computer vision, leveraging large amounts of pretraining data, conventional machine learning methods such as gradient boosted trees and their variants remain the predominant state-of-the-art (SOTA) across tabular prediction benchmarks. Recently, however, applying the in-context learning (ICL) paradigm to tabular prediction tasks has shown promising results by enabling pretraining of deep learning models across a large set of heterogeneous tables, constituting a new SOTA on small to medium tabular prediction tasks [17]. In this setting, predicting a target value $y$ based on the input features $\boldsymbol{x}$ of a row in a table $T$ additionally uses further rows from $T$ (the context), including their target values. This enables the model to adapt to new, unseen prediction problems at inference time, removing the need for task-specific fine-tuning.
@@
This approach was pioneered by the transformer-based TabPFN [16]. Pretrained on large amounts of synthetically generated classification tasks, its latest incarnation TabPFNv2 [17] produces SOTA results on tabular datasets with up to 10 0001000010\,00010 000 samples for both classification and regression tasks. In recent work, TabICL [28] extends the success story of this approach to even larger datasets, using special tabular embedding modules improving on the quadratic scaling in both the number of features and rows present in the TabPFN architecture.
@@
Common to both TabPFN and TabICL is that they are trained entirely on synthetically generated numerical data, with categorical features produced by an indexing procedure. While using synthetic data has many advantages, in particular its diversity at scale, a consequence is that the data does not contain any semantically meaningful values as found in real-world applications, both in the form of column names, and categorical or free-text entries. Furthermore, such data contains no additional data types such as times or dates that are abundant in practice. Consequently, these models do not utilize such information in a semantically meaningful way even when it is present at inference. In particular, column names are not used in either TabPFN or TabICL, and categorical features are encoded via one-hot or ordinal encoding disregarding any underlying semantics. We argue that semantic understanding can be captured by training on a large number of real-world tabular datasets. A primary example of this philosophy is TabuLa-8B [11], turning the pretrained large-language model (LLM) Llama 3-8B [7] into a tabular ICL model by fine-tuning it on around 3 million tables of the T4 dataset proposed therein. However, utilizing pretrained LLMs for tabular tasks has several limitations: Most importantly, textual serialization and tokenization of the input table is not token efficient, effectively limiting the maximum context length that can be processed. For example, TabuLa-8B operates on a maximum of 32 context rows. Furthermore, the tokenization schema and autoregressive nature of LLMs are not adapted to the tabular structure, resulting in a linear non-uniform token sequence, as cell values can be tokenized into different amounts of tokens, losing the 2D structure of the underlying data. Finally, the serialization and autoregressive processing result in an architecture that is neither row nor column permutation invariant – a property often desirable for tabular data [32].
@@
Aiming to bridge these approaches, we propose a table-native ICL model trained on the real-world T4 dataset [11], using embeddings tailored to different data modalities, in particular incorporating semantic embeddings of column names and categorical values. The resulting model is competitive to existing table-native ICL approaches across a range of tabular prediction benchmarks (OpenML-CC18 [1], OpenML-CTR23 [10], TALENT [37], and TabReD [30]), and achieves a new SOTA for the semantically rich CARTE benchmark [22], in particular in the low-data regime.
@@ 2 Related Work
Tabular deep learning:  Prediction on tabular data has traditionally been dominated by decision tree algorithms, particularly boosted variants like XGBoost [2], LightGBM [21], and CatBoost [27]. These models deliver strong performance but require separate training for each dataset and cannot leverage pretraining. Hence, much work has been done to transfer the success of deep learning methods and general pretraining to the tabular setting. Early examples include the FT-Transformer [13] and Xtab [39], whereas only more recent approaches have successfully shown consistently good performance overtaking boosted trees, for example TabR [12], RealMLP [18], or CARTE [22].
In-context learning on tabular data:  TabPFN [16] broke the long-standing dominance of boosted trees on small classification tasks, outperforming them by using row-level ICL. Pretrained on a large amount of synthetic tabular data, it generalizes to new tasks at inference time without fine-tuning or hyperparameter optimization. A recent variant, TabDPT [24], showed that equally excellent results can be achieved by training on real-world data using similarity-based retrieval for the context examples – an idea previously investigated in the TabR approach [12] – and further unlocked regression in this setting. Generalizing the row-based encoding, cell-based ICL introduced with TabPFNv2111In the following, we will focus solely on TabPFNv2 and refer to it simply as TabPFN. [17] and utilized also by TabICL [28] extended this success to larger datasets with up to 10 0001000010\,00010 000 and more samples, even outperforming SOTA AutoML solutions such as AutoGluon [9] on certain benchmarks.
Semantics and real data:  Capturing the rich semantics of real-world tabular datasets is a desirable property of a tabular foundation model, enabling the transfer of world knowledge across prediction tasks in addition to statistical patterns. The CARTE [22] architecture enables pretraining across a variety of real-world sources while capturing table semantics. It achieves SOTA results on the semantically rich CARTE benchmark, although it requires task-specific fine-tuning.
Modern LLMs have both deep semantic understanding and extensive world knowledge. Several works approach tabular ICL by tuning LLMs on tabular tasks, e.g. TabLLM [15], LIFT [6], or TabuLa-8B [11]. In particular, the works by Gardner et al. [11] are note-worthy for curating the T4 dataset, containing roughly 3 M tables extracted and processed from the TabLib collection [8] and for its excellent results in the very low data-regime.
@@ 3.1 Encoding
Text:  We transform each text cell to an embedding vector using a pretrained text embedding model. Note that we apply this to both free text columns as well as categorical columns, which can then retain the meaning in their labels. Any off-the-shelf embedding model can be used for this purpose – e.g. many are available based on the BERT architecture [5]. As many values have to be embedded for each table, there is a natural trade-off between accuracy and speed. We settle for a comparably small and fast model such as all-MiniLM-L6-v2 [19, 34], since the amount of semantic meaning it captures is already much richer than conventional categorical encodings, and we defer to Section 5.1 for an experimental ablation. Since the embeddings are potentially of a different dimension than the target embedding dimension $d$, we apply a learnable linear layer to the text embeddings.
Date:  In order to endow our model with knowledge of both the relative meaning of a dates (e.g. being able to compare two dates) and special dates (e.g. recurring holidays), we embed each of the numbers representing day, month, and year separately and sum the three resulting vectors. In particular, this is more token efficient then encoding dates into multiple features as is common in established preprocessors such as the one used by AutoGluon.
Numerical:  As numbers do not contain semantic meaning beyond their value, we apply a one-dimensional encoding. To make this procedure more robust during training, first, we clip columns between the 2% and 98% quantiles of the distribution. Second, we scale them to have zero mean and unit variance as is common. By Chebyshev’s inequality, this bounds resulting values to the interval $(-7.1,7.1)$, thus avoiding exploding gradients during training. Finally, the resulting number is multiplied by a learnable vector and a bias is added. If the original value was NaN, 0 is used instead, so the bias works as an “is-NaN” flag. An alternative embedding scheme is described in 3.4.
Column headers:  We embed column headers with the same model used for text cells. The result is passed through a separate learnable linear layer to map to the correct target dimension and summed with the cell embedding. Note that all the above embeddings are fully equivariant under permutations of either rows or columns – a property that is often desirable in the tabular domain. This equivariance makes the predictions more reproducible and robust and eliminates some of the need (such as the mapping of category to id, or of column to id) for bagging in models such as TabPFN. Semantic embeddings of cell values have been previously investigated in other works [35, 36, 22, 31], however, details of how the information is consumed differ in each implementation.
@@ 3.2 Backbone
We leave the TabPFN architecture mostly unchanged, with alternating “horizontal” (cross-column) and “vertical” (cross-row) self-attention transformer layers. In cross-column attention, each row is considered as a different batch element, and conversely. Following TabPFN, cross-column attention has no masking, while cross-row attention is masked so that each row can only attend to the provided context. To increase the modularity of our code, the feedforward MLP block of the transformer encoder is repeated after each self-attention, so that “horizontal” and “vertical” blocks have exactly the same weights. For sizing, we use the nomenclature popularized by BERT, e.g. by calling “base” the variant with 12 layers and a hidden dimension of 768. However, the real number of (non-embedding) weights is twice that of a BERT model, due to the presence of both “horizontal” and “vertical” layers.
Model weights can also be optionally shared between each instance of the transformer block, which consists of two interleaved attention layers. Such an architecture can be interpreted as a recurrent neural network (RNN), but unrolled in depth rather than in time. Such an iterative structure results in parameter efficiency and the possibility of stacking more blocks in comparison to the traditional approach. Empirically, we observed that sharing weights did not affect model performance and thus we use weight sharing as the default option for our model in the following.
@@ 3.4 Alternative architectures
Supervised clustering head:  We introduce an alternative method to perform classification that has the advantage of retaining semantic meaning not only for features but also for labels. As further benefits, full equivariance is preserved, without the need to map classes to IDs, and the limitation on the number of classes supported for prediction is lifted, improving upon the constraint currently present in many tabular ICL approaches such as TabPFN, TabDPT, or TabICL. For this, we take inspiration from Polewczyk and Spinaci [26]: For each row, the final output corresponding to the target column is mapped by a two-layer MLP to a vector $x$. We then build a matrix of shape $(n_{\mathrm{query}},n_{\mathrm{context}})$ by computing the cosine similarities between these vectors $x_{i}$, $x_{j}$ for each query row $i$ and context row $j$. During training, this is compared against the adjacency matrix with value 1 if two rows belong to the same class, and 0 otherwise: We compute the element-wise binary cross-entropy loss between this adjacency matrix’s entries and the clipped cosine similarities. That is, if $x_{i}$ is the vector from above for row $i$ and $c_{i}$ is its class, then the loss is given by
@@
Induced Set Attention Blocks:  One of the main limitations of the standard multi-head attention is its quadratic complexity. While memory can be effectively compressed to linear [29, 4], that is not the case for runtime. This poses a significant challenge when dealing with large tables. Taking inspiration from [23, 20, 28], we experiment with methods to handle larger context more efficiently.
Our proposed architecture replaces some of the multihead attention blocks within TabPFN with ISAB blocks from [23]. Following the notation from [23], let $\mathrm{MAB}(X,Y)$ denote the original multihead attention block from the transformer architecture, followed by a feedforward block including skip connections and layer normalization. Let $I$ be so-called inducing points [23], an additional set of parameters of shape $(n,d)$ for an arbitrary $n$, with $d$ denoting the hidden dimension, and define
@@
[…]
@@ 4.1 Training and inference
For pretraining, we use the T4 dataset [11]. We discard tables with fewer than 150 rows, which leaves $2.18$ M tables with a median of 750 rows and 9 columns. We randomly select 1000 rows, then between 50 and 900 rows as query, and use the rest as context. Then we randomly select one target column, excluding all date columns, numerical columns with more than 50% NaN values, and other columns having more than 20% of unique values. Finally, we up-sample non-numeric columns to have roughly the same proportion of regression and classification tasks. We train each model for between 4 and 10 million steps (i.e., 2 to 5 epochs) until convergence. We use a micro batch size of 1 and accumulate gradients to simulate a batch size of 256 (or 128 for smaller models of “mini” size). To improve stability, we employ gradient clipping and the AdamW optimizer with a maximum learning rate of $10^{-4}$, reached after a linear warm-up phase of 1000 gradient updates. Under this setup, we train a “base” sized model on a single H100 GPU, reaching a throughput of roughly 10 tables/s. Hence, full training takes between 4 and 12 days depending on the number of steps. Using the default parameters $n=12$, $d=768$, $d_{ff}=3072$ would result in a total of 172M parameters, but due to weight sharing there are only 16M trainable parameters.
We also experimented with curriculum learning, by further adding in a second step using the same training data as Ma et al. [24]. This data has fewer tables, 123, but many more cells in each table with a median of 11 k rows and 34 columns. In this second step, we increased the number of rows used for training to 4000. We refer to Section 5.1 for an analysis of the impact of training data size on model performance, which we found to be crucial.
At inference time, we apply 8-fold bagging, similar to the ensembling used by TabPFN. That is, from the original train split of a given evaluation dataset, we sample 8 times $c$ context rows with replacement, make a separate prediction with each collection of $c$ rows using the same model, and average the predictions (regression values or classification probabilities). While, during the default training, $c$ never surpasses 950, we can scale this during evaluation as much as memory permits. The combination with bagging allows ConTextTab to use up to $8c$ points as context. To limit runtime, in each bag, we also sample up to 500 columns if the original dataset has more than that.

### v4

@@ Abstract
Tabular in-context learning (ICL) has recently achieved state-of-the-art (SOTA) performance on several tabular prediction tasks. Previously restricted to classification problems on small tables, recent advances such as TabPFN [tabpfnv2] and TabICL [tabicl] have extended its use to larger datasets. Although current table-native ICL architectures are architecturally efficient and well-adapted to tabular data structures, their exclusive training on synthetic data limits their ability to fully leverage the rich semantics and world knowledge contained in real-world tabular data. At the other end of the spectrum, tabular ICL models based on pretrained large language models such as TabuLa-8B [tabula8b] integrate deep semantic understanding and world knowledge but are only able to make use of a small amount of context due to inherent architectural limitations. With the aim to combine the best of both these worlds, we introduce ConTextTab, integrating semantic understanding and alignment into a table-native ICL framework. By employing specialized embeddings for different data modalities and by training on large-scale real-world tabular data, our model is competitive with SOTA across a broad set of benchmarks while setting a new standard on the semantically rich CARTE benchmark. Code and model checkpoints are available at: https://github.com/SAP-samples/sap-rpt-1-oss.
**footnotetext: Equal contribution.
@@ 1 Introduction
Tables with information spread across rows and columns remain a predominant data format in many real-world applications [chui2018notes], making their understanding through machine learning algorithms critical. Despite the great success of deep learning approaches in natural language processing and computer vision, leveraging large amounts of pretraining data, conventional machine learning methods such as gradient boosted trees and their variants remain the predominant state-of-the-art (SOTA) across tabular prediction benchmarks. Recently, however, applying the in-context learning (ICL) paradigm to tabular prediction tasks has shown promising results by enabling pretraining of deep learning models across a large set of heterogeneous tables, constituting a new SOTA on small to medium tabular prediction tasks [tabpfnv2]. In this setting, predicting a target value $y$ based on the input features $\boldsymbol{x}$ of a row in a table $T$ additionally uses further rows from $T$ (the context), including their target values. This enables the model to adapt to new, unseen prediction problems at inference time, removing the need for task-specific fine-tuning.
@@
This approach was pioneered by the transformer-based TabPFN [tabpfnv1]. Pretrained on large amounts of synthetically generated classification tasks, its latest incarnation TabPFNv2 [tabpfnv2] produces SOTA results on tabular datasets with up to 10 00010\,000 samples for both classification and regression tasks. In recent work, TabICL [tabicl] extends the success story of this approach to even larger datasets, using special tabular embedding modules improving on the quadratic scaling in both the number of features and rows present in the TabPFN architecture.
@@
Common to both TabPFN and TabICL is that they are trained entirely on synthetically generated numerical data, with categorical features produced by an indexing procedure. While using synthetic data has many advantages, in particular its diversity at scale, a consequence is that the data does not contain any semantically meaningful values as found in real-world applications, both in the form of column names, and categorical or free-text entries. Furthermore, such data contains no additional data types such as times or dates that are abundant in practice. Consequently, these models do not utilize such information in a semantically meaningful way even when it is present at inference. In particular, column names are not used in either TabPFN or TabICL, and categorical features are encoded via one-hot or ordinal encoding disregarding any underlying semantics. We argue that semantic understanding can be captured by training on a large number of real-world tabular datasets. A primary example of this philosophy is TabuLa-8B [tabula8b], turning the pretrained large-language model (LLM) Llama 3-8B [dubey2024llamareducedauthors] into a tabular ICL model by fine-tuning it on around 3 million tables of the T4 dataset proposed therein. However, utilizing pretrained LLMs for tabular tasks has several limitations: Most importantly, textual serialization and tokenization of the input table is not token efficient, effectively limiting the maximum context length that can be processed. For example, TabuLa-8B operates on a maximum of 32 context rows. Furthermore, the tokenization schema and autoregressive nature of LLMs are not adapted to the tabular structure, resulting in a linear non-uniform token sequence, as cell values can be tokenized into different amounts of tokens, losing the 2D structure of the underlying data. Finally, the serialization and autoregressive processing result in an architecture that is neither row nor column permutation invariant – a property often desirable for tabular data [towardsfoundation].
@@
Aiming to bridge these approaches, we propose a table-native ICL model trained on the real-world T4 dataset [tabula8b], using embeddings tailored to different data modalities, in particular incorporating semantic embeddings of column names and categorical values. The resulting model is competitive to existing table-native ICL approaches across a range of tabular prediction benchmarks (OpenML-CC18 [openmlcc18], OpenML-CTR23 [openmlctr23], TALENT [talent], and TabReD [tabred]), and achieves a new SOTA for the semantically rich CARTE benchmark [CARTE], in particular in the low-data regime.
@@ 2 Related Work
Tabular deep learning:  Prediction on tabular data has traditionally been dominated by decision tree algorithms, particularly boosted variants like XGBoost [xgboost], LightGBM [lightgbm], and CatBoost [catboost]. These models deliver strong performance but require separate training for each dataset and cannot leverage pretraining. Hence, much work has been done to transfer the success of deep learning methods and general pretraining to the tabular setting. Early examples include the FT-Transformer [ft-transformer] and Xtab [xtab], whereas more recent approaches have shown consistently good performance overtaking boosted trees, for example TabR [tabr], RealMLP [realmlp], CARTE [CARTE], TabM [tabm], or ModernNCA [modernnca].
In-context learning on tabular data:  TabPFN [tabpfnv1] broke the long-standing dominance of boosted trees on small classification tasks, outperforming them by using row-level ICL. Pretrained on a large amount of synthetic tabular data, it generalizes to new tasks at inference time without fine-tuning or hyperparameter optimization. A recent variant, TabDPT [tabdpt], showed that equally excellent results can be achieved by training on real-world data using similarity-based retrieval for the context examples – an idea previously investigated in the TabR approach [tabr] – and further unlocked regression in this setting. Generalizing the row-based encoding, cell-based ICL introduced with TabPFNv2111In the following, we will focus solely on TabPFNv2 and refer to it simply as TabPFN. [tabpfnv2] and utilized also by TabICL [tabicl] extended this success to larger datasets with up to 10 00010\,000 and more samples, even outperforming SOTA AutoML solutions such as AutoGluon [autogluon] on certain benchmarks.
Semantics and real data:  Capturing the rich semantics of real-world tabular datasets is a desirable property of a tabular foundation model, enabling the transfer of world knowledge across prediction tasks in addition to statistical patterns. The CARTE [CARTE] architecture enables pretraining across a variety of real-world sources while capturing table semantics. It achieves SOTA results on the semantically rich CARTE benchmark, although it requires task-specific fine-tuning.
Modern LLMs have both deep semantic understanding and extensive world knowledge. Several works approach tabular ICL by tuning LLMs on tabular tasks, e.g. TabLLM [tabllm], LIFT [lift], or TabuLa-8B [tabula8b]. In particular, the works by tabula8b are note-worthy for curating the T4 dataset, containing roughly 3 M tables extracted and processed from the TabLib collection [tablib] and for its excellent results in the very low data-regime.
@@ 3.1 Encoding
Text:  We transform each text cell to an embedding vector using a pretrained text embedding model. Note that we apply this to both free text columns as well as categorical columns, which can then retain the meaning in their labels. Any off-the-shelf embedding model can be used for this purpose – e.g. many are available based on the BERT architecture [devlin-etal-2019-bert]. As many values have to be embedded for each table, there is a natural trade-off between accuracy and speed. We settle for a comparably small and fast model such as all-MiniLM-L6-v2 [all-MiniLM-L6-v2, wang2020minilm], since the amount of semantic meaning it captures is already much richer than conventional categorical encodings, and we defer to Section LABEL:sec:ablation for an experimental ablation. Since the embeddings are potentially of a different dimension than the target embedding dimension $d$, we apply a learnable linear layer to the text embeddings.
Date:  In order to endow our model with knowledge of both the relative meaning of a dates (e.g. being able to compare two dates) and special dates (e.g. recurring holidays), we embed each of the numbers representing day, month, and year separately and sum the three resulting vectors. In particular, this is more token efficient than encoding dates into multiple features as is common in established preprocessors such as the one used by AutoGluon.
Numerical:  As numbers do not contain semantic meaning beyond their value, we apply a one-dimensional encoding. To make this procedure more robust during training, first, we clip columns between the 2% and 98% quantiles of the distribution. Second, we scale them to have zero mean and unit variance as is common. By Chebyshev’s inequality, this bounds resulting values to the interval (7.1,7.1)(-7.1,7.1), thus avoiding exploding gradients during training. Finally, the resulting number is multiplied by a learnable vector and a bias is added. If the original value was NaN, 0 is used instead, so the bias works as an “is-NaN” flag. An alternative embedding scheme is described in 3.4.
Column headers:  We embed column headers with the same model used for text cells. The result is passed through a separate learnable linear layer to map to the correct target dimension and summed with the cell embedding.
After summation, the embeddings are normalized via layer normalization. Note that all the above embeddings are fully equivariant under permutations of either rows or columns. This equivariance makes the predictions more reproducible and robust and eliminates some of the need (such as the mapping of category to ID, or of column to ID) for bagging in models such as TabPFN. Semantic embeddings of cell values have been previously investigated in other works [yak2023ingestables, ye2024crosstable, CARTE, PORTAL], however, details of how the information is consumed differ in each implementation.
@@ 3.2 Backbone
We leave the TabPFN architecture mostly unchanged, with alternating “horizontal” (cross-column) and “vertical” (cross-row) self-attention transformer layers. In cross-column attention, each row is considered as a different batch element, and conversely. Following TabPFN, cross-column attention has no masking, while cross-row attention is masked so that each row can only attend to the provided context. To increase the modularity of our code, the feedforward MLP block of the transformer encoder is repeated after each self-attention block so that “horizontal” and “vertical” blocks have the same structure. For sizing, we use the nomenclature popularized by BERT, e.g. calling “base” the variant with 12 layers and a hidden dimension of 768. However, the real number of (non-embedding) weights is twice that of a BERT model, due to the presence of both “horizontal” and “vertical” layers.
Model weights can be optionally shared between each instance of the transformer block, consisting of two interleaved attention layers. Such an architecture can be interpreted as a recurrent neural network, unrolled in depth rather than in time. Such an iterative structure results in parameter efficiency and the possibility of stacking more blocks in comparison to the traditional approach. Empirically, we observed that sharing weights did not affect model performance and thus we use weight sharing as the default option for our model in the following.
@@ 3.4 Alternative architectures
Supervised clustering head:  We introduce an alternative method to perform classification that has the advantage of retaining semantic meaning not only for features but also for labels. As further benefits, full equivariance is preserved, without the need to map classes to IDs, and the limitation on the number of classes supported for prediction is lifted, improving upon the constraint currently present in many tabular ICL approaches such as TabPFN, TabDPT, or TabICL. For this, we take inspiration from clustertabnet: For each row, the final output corresponding to the target column is mapped by a two-layer MLP to a vector $x$. We then build a matrix of shape $(n_{\mathrm{query}},n_{\mathrm{context}})$ by computing the cosine similarities between these vectors $x_{i}$, $x_{j}$ for each query row $i$ and context row $j$. During training, this is compared against the adjacency matrix with value 1 if two rows belong to the same class, and 0 otherwise: We compute the element-wise binary cross-entropy loss between this adjacency matrix’s entries and the clipped cosine similarities. That is, if $x_{i}$ is the vector from above for row $i$ and $c_{i}$ is its class, then the loss is given by
@@
Induced Set Attention Blocks:  One of the main limitations of the standard multi-head attention is its quadratic complexity. While memory can be effectively compressed to linear [efficientattention, flashattention], that is not the case for runtime. This poses a significant challenge when dealing with large tables. Taking inspiration from [isab, perceiver, tabicl], we experiment with methods to handle larger context more efficiently.
Our proposed architecture replaces some of the multihead attention blocks within TabPFN with ISAB blocks from [isab]. Adopting the notation from [isab], let $\mathrm{MAB}(X,Y)$ denote the standard multihead attention block from the transformer architecture, where $X$ represents the queries and $Y$ serves as both the keys and values. This is followed by a feedforward block, a skip connection, layer normalization, and ultimately an intermediate block. Let $I$ be so-called inducing points [isab], an additional set of parameters of shape $(n,d)$ for an arbitrary $n$, with $d$ denoting the hidden dimension, and define
@@
where $C$ and $Q$ denote context and query rows, respectively, and $\mathbin{+\mkern-10.0mu+}$ denotes concatenation. In each attention computation, we apply masking so that only context and inducing vectors are attended to. For a detailed overview of the architecture, refer to Appendix Figure LABEL:fig:ISAB_diagram.
@@ 4.1 Training and inference
For pretraining, we use the T4 dataset [tabula8b]. We discard tables with fewer than 150 rows, which leaves $2.18$ M tables with a median of 750 rows and 9 columns. We randomly select 1000 rows, then between 50 and 900 rows as query, and use the rest as context. Subsequently, we randomly select one target column, excluding all date columns, numerical columns with more than 50% NaN values, and other columns having more than 20% of unique values. Finally, we up-sample non-numeric columns to have roughly the same proportion of regression and classification tasks. We train each model for between 4 and 10 million steps (i.e., 2 to 5 epochs) until convergence. We use a micro batch size of 1 and accumulate gradients to simulate a batch size of 256 (or 128 for smaller models of “mini” size). To improve stability, we employ gradient clipping and the AdamW optimizer with a maximum learning rate of $10^{-4}$, reached after a linear warm-up phase of 1000 gradient updates. Under this setup, we train a “base” sized model on a single H100 GPU, reaching a throughput of roughly 10 tables/s. Hence, full training takes between 4 and 12 days depending on the number of steps. Using the default parameters $n=12$, $d=768$, $d_{\textrm{ff}}=3072$ results in a total of 172 M parameters which is reduced to 16 M trainable parameters when using weight sharing.
We also experimented with curriculum learning, by further adding in a second step using the same training data as tabdpt. This data has fewer tables, 123, but many more cells in each table with a median of 11 k rows and 34 columns. In this second step, we increased the number of rows used for training to 4000. We refer to Section LABEL:sec:ablation for an analysis of the impact of training data size on model performance, which we found to be crucial.
At inference, we apply 8-fold bagging, similar to the ensembling used by TabPFN. That is, from the original train split of a given evaluation dataset, we sample 8 times $c$ context rows with replacement, make a separate prediction with each collection using the same model, and average the predictions (regression values or classification probabilities). While, during the default training, $c$ never surpasses 950, we can scale this during evaluation as much as memory permits. The combination with bagging allows ConTextTab to use up to $8c$ points as context. By default, we use a context size of $c=8192$ at inference. To limit runtime, in each bag, we sample up to 500 columns if the dataset has more.